# Accounting for multiplicity in machine learning benchmark performance

### v4

@@ 3.3 The probability distribution of $\hat{\theta}^{\,max}_{SOTA}$
Let $n$ denote the size of the test set, $m$ be the number of classifiers, and $\theta$ be the probability of correct prediction. Let $X$ be the random variable that denotes the number of failures (wrong predictions) for a given classifier. We have that $X\sim\mathcal{B}(n,1\text{-}\theta)$. We denote by $P_{x}$ the cumulative distribution $P(X\leq x)$. The random sample, $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ consists of the number of wrong predictions for each classifier $j=1,\ldots,m$ on a test set. Let $C_{x}$ be the random variable that denotes the number of classifiers with at most $x$ failures. We have that $C_{x}\sim\mathcal{B}(m,P_{x})$, because each classifier will independently have a probability $P_{x}$ of having at most $x$ wrong predictions.
@@
We are interested in the probability of at least one classifier having at most $x$ wrong predictions, because it corresponds to the probability function of the $\hat{\theta}^{\,max}_{SOTA}$ estimator, $P(\hat{\theta}^{\,max}_{SOTA}>\frac{n-x}{n})$, in other words, at least one $\hat{\theta}_{j}>\frac{n-x}{n}$. We can derive this from $C_{x}$:
$P(C_{x}>0|m,n,\theta).$ (1)
Note that this corresponds to the cumulative distribution function (cdf) $F(z)$ for the random variable $Z$, denoting the most wrong predictions among all classifier. We have cdf $F(z)$, the probability that the most wrong predictions among all classifiers is less than or equal to a number $z$, can easily be derived from the binomial distribution
@@
Let $f(z)$ denote the probability mass function of $Z$. We then have
@@
The last line is derived by binomial expansion. The practical relevance of the binomial expansion is limited by the need to approximate binomial coefficient for large $m$. From Eq. 3.3, the expected value and variance can easily be calculated. An example of $F(z)$ can be seen in Fig. 2, and the corresponding $f(z)$ in Fig. 3. The pmf $f(z)=P(Z=z)$ denotes the probability of at least one classifier having exactly $z$ wrong predictions, and no classifiers having fewer than $z$ wrong predictions.
@@ 3.4 A simulated public competition example
The multiplicity mechanisms described above apply to public competitions where several teams submit their predictions on a test set. Winning teams are those with highest accuracy, $\max_{j}\hat{\theta}_{j}$, where $\hat{\theta}_{j}=(n-x_{j})/n$ is the accuracy of a classifier $j$ on the test set. This is often referred to as the state of the art, and hence, we have that $\hat{\theta}^{\,max}_{SOTA}(\mathbf{x})=\max_{j}\hat{\theta}_{j}$. As exemplified in Section 3.2, multiplicity biases the estimation of $\theta_{SOTA}$, and Section 3.3 gives us the tools to analyze its impact. In the following, we will use a significance level of $\alpha=0.05$ and two-sided confidence intervals.
@@
Consider a classification problem with a test set of size $n=3,000$, and a classifier with probability of correct prediction $\theta=0.90$. The pmf of a single classifier is shown as a histogram in Fig. 1. The 95% confidence interval for $\theta$ if we have observed $\hat{\theta}=0.90$ is $(0.8887,0.9105)$, and we denote it by $(\theta_{\alpha/2},\theta_{1\text{-}\alpha/2})$.
@@
If there are $m=1,000$ classifiers, each with probability of correct prediction $\theta=0.90$, the probability of at least one team achieving at least $\theta_{1\text{-}\alpha/2}$ is close to 1. This comes as no surprise given the definition of a confidence interval. It means that we can be almost certain that the accuracy of the top ranked classifier, which is the default SOTA, is above the upper limit of the confidence interval of the true performance.
On the other hand, it is straightforward to show that if each individual classifier has a probability of $P_{x}=0.00002532$ for at most $x$ failures, then, with a probability of $\alpha/2$, at least one team will achieve at least $\hat{\theta}=0.9213$. We denote this by $\theta^{m}_{1\text{-}\alpha/2}$, the multiplicity-adjusted upper limit of the $(1\text{-}\alpha)$ confidence interval.
Similarly, the expected value of $\hat{\theta}^{\,max}_{SOTA}$ can easily be calculated from Eq. 3.3, and we have $E(\hat{\theta}^{\,max}_{SOTA})=0.9173$. Both are optimistic estimates, knowing that $\theta_{SOTA}$ is $0.90$.
@@
Within the competition context, this is not necessarily a problem, because each team has the same probability of achieving the highest accuracy given that they have the same $\theta$. In a scientific context, using $\max_{j}\hat{\theta}_{j}$ as an estimate for $\theta_{SOTA}$ without multiplicity corrections is problematic. Consider a new classifier with probability of correct prediction $\theta_{m+1}=\theta_{1\text{-}\alpha/2}$ from above, displayed as the red curve in Fig. 1. By common scientific standards, the expected performance, $E(\hat{\theta}_{m+1})=\theta_{m+1}$ is considered significantly better than the expected performance of any of the $1,000$ classifiers with $\theta=0.90$, and is of interest to the research community. But the probability of this classifier beating the upper limit of the multiplicity adjusted confidence interval is very low: $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})=0.0190$, and can be seen as the red area in Fig. 1. Similarly, $P(\hat{\theta}_{m+1}\geq E(\hat{\theta}^{\,max}_{SOTA}))=0.0996$, seen as the blue plus red area in Fig. 1. This means that by using the maximum accuracy as SOTA, we can expect that 9 out of 10 classifiers that are in fact better will still be naively considered worse.
The probability mass functions of single $\hat{\theta}$.
Figure 1: The grey histogram shows the pmf of a classifier with probability of correct prediction $\theta$. The red curve is the pmf of a classifier with probability of correct prediction $\theta_{1\text{-}\alpha/2}$, the upper limit of the $95\%$ CI for $\hat{\theta}(\mathbf{x})=\theta$. $E(\hat{\theta}^{\,max}_{SOTA})$ denotes the expected value of the highest ranked among $m=1,000$ classifiers. $\theta^{m}_{1\text{-}\alpha/2}$ is the multiplicity corrected upper limit of the $95\%$ CI. The shaded areas corresponds to $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})$ (red) and $P\big{(}\hat{\theta}_{m+1}\geq E(\hat{\theta}^{\,max}_{SOTA})\big{)}$ (blue+red).
@@
The cumulative distribution function, $F(z)$ from Eq. 3.3, is displayed in Fig. 2, and describes the cumulative probability of at least one classifier having at most $z$ failures. Or, in other words, the cumulative probability of maximum accuracy. The number of failures are displayed on the top horizontal axis, and the corresponding accuracy on the bottom horizontal axis. Note that the horizontal axis is reversed compared to Fig. 1.
Cumulative distribution function, $F(z|m,n,\theta)$, for at least one classifier having at most $z$ failures.
Figure 2: The figure shows the cumulative distribution function for at least one classifier having at most $z$ failures. The horizontal axis denotes the number of failures, $z$, on top and the accuracy at the bottom.
The corresponding pmf, $f(z)$ from Eq. 3.3, is displayed as a histogram in Fig. 3.
Probability mass function, $f(z|m,n,\theta)$, for the smallest number of failures, $z$, among all classifiers.
Figure 3: The figure shows the corresponding pmf for Fig. 2. The horizontal axis denotes the number of failures, $z$, on top and the corresponding accuracy at the bottom.
From Eq. 3.3 we get the expected value and variance. For $m=1,000$, $n=3,000$ and $\theta=0.9$ this corresponds to $E(\hat{\theta}^{\,max}_{SOTA})=0.9173$ with a standard deviation of $0.001817$. We can see how different values for $m$, $n$ and $\theta$ affects the $\hat{\theta}^{\,max}_{SOTA}$ distribution in Table 1.
Expected values and standard deviations for $\hat{\theta}^{\,max}_{SOTA}$
m n $\theta$ $E(\hat{\theta}^{\,max}_{SOTA})$ $\sigma_{\hat{\theta}^{\,max}_{SOTA}}$
1,000 3,000 0.85 0.8707 0.002197
@@
Table 1: The expectation and standard deviation for the SOTA estimator with varying test set size, number of classifiers and probability of correct prediction. The bold line is the example used in Fig. 3.
4 Non-i.i.d.
The probability distribution in Section 3.4 is based on independent classifiers with identical $\theta$s. This assumption is rarely true in practice. A classifier’s ability to perform better than random guesses is due to correlation between features and labels in a training set. The teams will pick up on many of the same features and correlations, regardless of which method they choose. In addition, many teams will use different versions of related methods. We investigate dependency and non-identical classifiers, and the impact on the bias. Table 2 gives an overview, details are in the following sections.
Upper bounds, expected values and standard deviations for $\hat{\theta}^{\,max}_{SOTA}$
section independent identical upper 95% CI expected value standard deviation
3.4 YES YES 0.9213 0.9173 0.001817
4.1 NO YES 0.9207 0.9140 0.003481
4.1 NO YES 0.9173 0.9140 0.001484
4.2 YES NO 0.9177 0.9130 0.002129
4.3 NO NO 0.9173 0.9101 0.003649
@@
Table 2: The upper bound of the 95% CI, the expectation and standard deviation for the SOTA estimator for different combinations of independency and identicality. The true state-of-the-art is $\theta=0.90$.
4.1 Dependent, identical classifiers
Let $Y_{j}$ be the Bernoulli variable with value $1$ if classifier $j$ predicts a data point correctly. We have $P(Y_{j}=1)=\theta_{j}$, and $\theta_{j}=\theta$ for classifiers with equal probability of success. Mania et al. (2019) calculated the similarity, defined as the fraction of two models giving the same output, which is equivalent to
$P\big{(}\{Y_{j},Y_{k}\}=\{0,0\}\big{)}+P(\{Y_{j},Y_{k}\}=\{1,1\})$ (4)
$=(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}$
for independent models. For dependent models with correlation $\rho_{jk}$, this becomes
$(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}+2\rho_{jk}\sqrt{\theta_{j}\theta_{k}(1-\theta_{j})(1-\theta_{k})}.$ (5)
By setting $\theta_{j}=\theta$ and $\rho_{jk}=\rho$, $\rho$ can easily be estimated from the estimations of $\theta$, $P\big{(}\{Y_{j},Y_{k}\}=\{0,0\}\big{)}$ and $P(\{Y_{j},Y_{k}\}=\{1,1\})$. In Mania et al. (2019), they used the average accuracy for top-performing ImageNet models, $\hat{\theta}=0.756$. For details, see Mania et al. (2019). Note that these are multi-class problems, so a lower accuracy than for binary classification problems is to be expected. From their calculated similarity of $0.85$, the estimated correlation is $0.36$.
Boland et al. (1989) describes dependency in the majority systems context, with $\rho_{0j}=\text{corr}(Y_{0},Y_{j}),\,j=1,\ldots m$, and the $Y_{j}$s are independent given $Y_{0}$. This gives $\rho_{jk}=\rho_{0j}^{2}$. The observations from $Y_{0}$ does not contribute directly to $\hat{\theta}^{\,max}_{SOTA}$, but can be seen as, e.g., the influence that a common training set has on the classifiers.
The binomial distribution for dependent Bernoulli trials is not an easy problem, see, e.g., Ladd (1975) and Hisakado et al. (2006). We here demonstrate by simulation how dependency reduces the bias. We use the estimated correlation from Mania et al. (2019), which corresponds to $\rho_{0j}=0.6$ in the multiplicity set-up with an external $Y_{0}$.
We use the set-up from Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$, and probability of correct prediction $\theta=0.90$. We simulate two scenarios; one where $Y_{0}$ is random, and one where the $y_{0}$’s are fixed with exactly $(n-x)/n=0.90$ correct predictions. The fixed scenario can correspond to fixed training conditions, e.g., models pre-trained on ImageNet. The random scenario picks up the randomness in, e.g., a training set. Although all teams have the same training set, the training set is a random sample.
The upper bound of the $95\%$ confidence interval for the dependent classifiers is at $0.9173$ for the fixed $y_{0}$ and $0.9207$ for the random $Y_{0}$, but they have the same expected value.
Minimum number of failures in $m=1,000$ classifiers.
(a) $\rho_{0j}=0.6$, random $Y$
(b) $\rho_{0j}=0.6$, fixed $y$
Figure 4: The minimum number of failures in $100,000$ simulations, with probability of success $\theta=0.9$.
Note that, with $\theta=0.90$, the similarity as described by Mania et al. (2019) is $0.82$ for two independent classifiers. When the probability of correct prediction is high, there is little room for variation between classifiers, and the inherent similarity is high.

### v5

@@ 3.3 The probability distribution of $\hat{\theta}^{\,\max}_{SOTA}(\mathbf{X})$
Let $X$ be the random variable that denotes the number of failures (wrong predictions) for a given classifier, with $X\sim\mathcal{B}(n,1\text{-}\theta)$. The random sample, $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ consists of the number of wrong predictions for each classifier $j=1,\ldots,m$ on a sample. Let $C_{x}$ be the random variable that denotes the number of classifiers with at most $x$ failures. We have that $C_{x}\sim\mathcal{B}(m,P_{x})$.
@@
We are interested in the probability of at least one classifier having at most $x$ wrong predictions, because it corresponds to the probability function of the $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ estimator, $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}}\geq\frac{n-x}{n})$. We can derive this from $P(C_{x}>0|m,n,\theta)$. This corresponds to the cdf for the random variable $Z$ denoting the most wrong predictions among all classifiers:
@@
Let $f(z)=P(Z=z)$ denote the pmf of $Z$, the probability of at least one classifier having exactly $z$ wrong predictions, and no classifiers having fewer than $z$ wrong predictions. We then have
@@
The variable $Z$ corresponds to $\hat{\theta}_{SOTA}^{\,max}$ in the sense that the smallest number of failures among all classifiers corresponds to the sample maximum probability of success. Equations 3.3 and 3.3 can be used to analyse $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for a given situation, as we do in the next section, where the impact that the three parameters have on the bias is also presented.
@@ 3.4 A simulated public competition example
The multiplicity mechanism that we described in the coin-flip example applies to public competitions where multiple teams submit the predictions made by their classifiers on a test set. The classifier with highest performance is declared winner, and is often referred to as SOTA later on. As exemplified in Section 3.2, multiplicity biases the naive sample maximum $\theta_{SOTA}$-estimator, and Section 3.3 gives us the tools to analyze its impact. In the following, we will use a significance level of $\alpha=0.05$ and two-sided CIs.
@@
Consider a classification problem with a test set of size $n=3,000$, and a classifier with probability of success $\theta=0.90$. The pmf of a single $\hat{\theta}(X)$ is shown as the blue curve in Fig. 1, where the dotted line indicates $E(\hat{\theta})=\theta$. The 95% CI for $\hat{\theta}=0.90$ is $(0.8887,0.9105)$, and we denote it by $(\theta_{\alpha/2},\theta_{1\text{-}\alpha/2})$, displayed as a blue line on the horizontal axis.
@@
If there are $m=1,000$ independent classifiers, each with probability of success $\theta=0.90$, the probability of at least one classifier achieving at least $\hat{\theta}=\theta_{1\text{-}\alpha/2}$ is close to 1. This comes as no surprise given the definition of a CI. It means that we can be pretty certain that the sample performance of the top ranked classifier is above the upper limit of the CI of the expected performance.
@@
Consider a new classifier with probability of success $\theta_{m+1}=\theta_{1\text{-}\alpha/2}$. The corresponding pmf is displayed as the red curve in Fig. 1, with the dotted line indicating $E(\hat{\theta}_{m+1})=\theta_{1\text{-}\alpha/2}$. By common scientific standards, the expected performance is considered significantly better than the expected performance of any of the $1,000$ classifiers with $\theta=0.90$, and is of interest to the research community. However, the multiplicity effect can conceal this. In the competition context, this is not necessarily a problem, since each classifier has the same probability of achieving the highest sample performance given that they have the same $\theta$. In a scientific context, using $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as an estimate for $\theta_{SOTA}$ without multiplicity corrections is problematic.
The upper limit of the $(1\text{-}\alpha)$ CI of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ is $0.9213$, displayed on the horizontal axis in Fig. 1, denoted as $\theta^{m}_{1\text{-}\alpha/2}$. We have $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})=0.9173$, calculated from Eq. 3.3, also displayed on the horizontal axis in Fig. 1. The probability of the new classifier beating the upper limit of the sample maximum CI is very low: $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})=0.02$, illustrated by the darker green area in Fig. 1. Similarly, $P(\hat{\theta}_{m+1}\geq E(\operatorname{\hat{\theta}^{\,max}_{SOTA}}))=0.10$ seen as the entire green area in Fig. 1. This means that by using the sample maximum as an estimator for SOTA, we can expect that 9 out of 10 classifiers that are significantly better will still be naively considered worse.
The pmfs of two single $\hat{\theta}(X)$ compared to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ statistics
Figure 1: Against a mass of identical classifiers, a single significantly better classifier has little chance of beating the state of the art: The blue and the red curves show the pmfs of $\hat{\theta}(X)$ for classifiers with probability of success $\theta$ and $\theta_{1\text{-}\alpha/2}$, respectively. $\theta_{1\text{-}\alpha/2}$ is the upper limit of the CI of $\hat{\theta}(\mathbf{x})=\theta$. The CI is displayed as the blue horizontal line. $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ and $\theta^{m}_{1\text{-}\alpha/2}$ denote the expected value and the upper limit of the CI of the sample maximum estimator for $m=1,000$ classifiers. The shaded areas corresponds to $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})$ and $P\big{(}\hat{\theta}_{m+1}\geq E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})\big{)}$, and reflects the slim chances that the single better classifier has of beating the multiple worse classifiers.
For the entire distribution of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, see Fig. 2.
Cdf and pmf for at least one classifier having at most $z$ failures, corresponding to the sample maximum estimator
(a) Cdf $F(z)$, see Eq. 3.3
(b) Pmf $f(z)$, see Eq. 3.3
@@
Figure 2: The distribution and the probability function of the sample maximum estimator for $n=3,000$, $\theta=0.90$, and $m=1,000$. The horizontal axes denotes the number of failures, $z$, at the top and the corresponding sample maximum estimates, $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, at the bottom. Note that the horizontal axis is reversed compared to Fig. 1.
The distribution of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ will vary with $m$, $n$ and $\theta$, and Table 1 shows the expectation and standard deviation for a small collection of examples. To complement the table, Fig. 3 a)-c) show the bias of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as a function of $m$, $n$ and $\theta$, respectively.
Expected values and standard deviations for $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$
m n $\theta$ $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ $\sigma_{\operatorname{\hat{\theta}^{\,max}_{SOTA}}}$
@@
Table 1: The expected value and standard deviation for the SOTA estimator with varying test set size, number of classifiers and probability of correct success. The bold line is the example used in Fig. 2 and corresponds to the intersection between the gray vertical lines and the solid lines in Fig. 3.
For each curve, the two other parameters are kept constant, and each figure displays five curves, corresponding to Table 1.
The bias of the sample maximum estimator
Figure 3: The bias of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as a function of each of its three parameters; number of classifiers, test set size and probability of success. The intersection between the black solid lines and the gray vertical lines correspond to the bold line in Table 2. The dotted and dashed lines correspond to variation of parameter values found in Table 2, colour-coded accordingly.
In Fig. 3(a), we see that the impact of number of classifiers is dramatic, and multiplicity-considerations must be treated with full attention starting at $m=2$. We can see this from the cdf in Eq. 3.3, where $m$ is the exponent. In Fig. 3(b) is that the test set size also has large impact on the bias. With small test sets, the variance of each individual $\hat{\theta}_{j}(X)$ is high, increasing the probability of sample estimates far from $\theta$. As seen in Fig. 3(c), the bias decreases as $\theta$ approaches 1. The variance of the individual $\hat{\theta}_{j}(X)$s is at its highest for $\theta=0.5$ and decreases monotonically, and it is this variance that is reflected in the bias, similar to the effect of the test set size.
The change in absolute numbers may seem modest, but the practical implications are large, as demonstrated in Section 5, where we analyse the results from three public competitions.
4 Dependency and non-identical $\theta$s
The probability distribution in Section 3.4 is that of independent classifiers with identical $\theta$s, an assumption that is rarely true in practice. We expand our investigation to non-identical $\theta$s and dependency. Table 2 and Figure 5 give an overview of some exemplifying modelss, details are in the following sections.
Characteristics of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for dependent and non-identical classifiers
Section Independent Identical Upper 95% CI $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ $\sigma_{\operatorname{\hat{\theta}^{\,max}_{SOTA}}}$
3.4 YES YES 0.9213 0.9173 0.0018
4.1 YES NO 0.9177 0.9130 0.0021
4.2 NO YES 0.9207 0.9140 0.0035
4.3 NO NO 0.9173 0.9101 0.0036
@@ 4 Dependency and non-identical $\theta$s
Table 2: The upper bound of the 95% CI, the expected value and standard deviation for the sample maximum estimator for different combinations of independency and identicality. The true SOTA is $\theta=0.90$, the test set size is $n=3,000$, the number of classifiers is $m=1,000$. The correlation coefficient and the $\Theta$-vector are found in their respective subsections.
Bias for non-identical and dependent classifiers.
(a) Bias as a function of $\min(\mathbf{\Theta})$, with constant $\max(\mathbf{\Theta})=0.90$
(b) Bias as a function of correlation coefficient.
Figure 4: Bias of the sample maximum estimator as a function of $\min(\Theta)$ for non-identical classifiers and of $\rho_{0}$ for conditional independent classifiers. (a) The right tip of the curve is identical classifiers, and following the curve from right to left the $\mathbf{\Theta}$-vector stretches out, and the bias decreases. The intersection with the solid gray vertical line corresponds to the cdf and pmf in Fig. 5, and produces the CI, the expectation and the standard deviation in line two in Table 2. The intersections between the gray vertical lines correspond to the left tips of the black lines in Fig. 4(b), where the classifiers are independent. (b) The red line are identical classifiers, as discussed in Section 4.2. The left tip of the red curve corresponds to the i.i.d. classifiers in Section 3.4, and the intersection with the grey vertical line corresponds to line three in Table 2. The black lines show bias as a function of correlation coefficient demonstrated for different $\mathbf{\Theta}$-vectors. The intersection between the solid black curve and the grey vertical line corresponds to line four in Table 2. The scale of the $y$-axis is kept the same as in Fig. 3 for easier comparison of which parameters have the most impact.
Overall, Table 2 demonstrates that dependency and non-identical $\theta$s reduces the bias of the $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ estimator, and assumptions regarding dependency and identicality are crucial to include in a model. Figure 4 shows how the bias changes as a function of $\Theta$ and $\rho_{0}$ for these specific models.
We use the set-up from Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$, $\theta_{SOTA}=0.90$, and $100,000$ repetitions when simulations are needed. We refer to classifiers with identical $\theta$s as identical classifiers.